Twice the difference of a number and 8 is equal to three times the sum of the number and 3. Find the number.
1
Expert's answer
2015-02-12T02:51:56-0500
let number will be x
Denote twice the difference between the number and 8 as 2(x-8)
than 3 times sum of number and 3 will be 3(x+3)
So we get an equation
2(x-8)=3(x+3)
2x-16=3x+9
rearrange the terms and numbers
3x-2x= -16-9
x= -25